# Still More Shades of Null: An Evaluation Suite for Responsible Missing Value Imputation

### v4

@@ 4. Single and Multi-Mechanism Missingness
Different models are the best-performing on different datasets. In Figures 34 and 5, we report on the best-performing models (according to F1) for each dataset, and compare performance against a model trained on the clean training set using the same model architecture (performance shown as a blue line in the plots).
@@
Figure 3. F1 of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Datasets are in increasing order by size. Blue line shows performance of the model trained on clean data.
@@ 4.1. Correctness of the Predictive Model
Note that the F1 score on the bank dataset is extremely low (0.32). This is due to severe class imbalance in the dataset (the base rate is 0.117, as reported in Table 9 in Appendix B). Further, from Figure 20 in Appendix C.2, we can see that the accuracy of the lgbm__\__clf is 0.89, which is marginally better than the accuracy reported using Logistic Regression (0.885) in (Le Quy et al., 2022) in Table 15.
@@
Figure 4. True Positive Rate Difference (unfairness) of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets ($x$-axis), and missingness mechanisms (subplots). Values close to 0 are desirable. Datasets are in increasing order by size. Blue line shows TPRD of the model trained on clean data.
@@ 4.2. Fairness of the Predictive Model
Figure 5. Label Stability of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 1 are desirable. Datasets are ordered in increasing order by size. The blue line indicates performance of the model trained on clean data.
@@ 4.3. Stability of the Predictive Model
Figure 5 shows label stability of models trained using different MVI techniques. In line with conventional wisdom (Domingos, 2000), stability depends most strongly on dataset characteristics (mainly size) and model type. Impute-then-classify approaches like miss-forest, auto-ml and datawig, which were the best-performing according to F1, also match the stability of the model trained on clean data under all missingness types and dataset sizes. nomi, which performs best on accuracy and training time, demonstrates comparable stability to these top techniques on all datasets. clustering, which as was poor-performing according to F1 (correctness) is also poor-performing according to stability on small datasets (diabetes and german, with 905 and 1k samples, respectively). A notable and surprising exception is under mixed missingness on german, where clustering is competitive with other impute-then-classify approaches, despite not being competitive on MCAR, MAR or MNAR individually. The reason for this is that imputation accuracy of MVI techniques impacts data uncertainty, which then directly impacts overall model uncertainty (Gal et al., 2016).
Deletion worsens stability compared to the model trained on clean data for all missingness types on diabetes, but, notably, only under MNAR on german. boostclean, which constitutes its own model type, shows a deviation from the stability of the model trained on clean data on all datasets except bank and all missingness types: worsening stability compared to the clean model on folk-income, law-school and folk-employment (except under MCAR), but, surprisingly, improving it on heart, even under mixed missingness.
@@ 4.4. Imputation Quality and Fairness
Imputation Quality.
We find that MVI techniques with vastly different imputation quality can give rise to models with comparable F1 scores, indicating that imputation correctness is not sufficiently predictive of model correctness (Zhang and Long, 2021a). For example, in Figure 6(a) on the diabetes dataset, median-dummy has an F1 score (imputation) close to 0; auto-ml, miss-forest and nomi have an F1 score (imputation) close to 1; and all other imputers have an F1 score (imputation) between 0.5-0.6, but the models trained with all of these techniques have an F1 close to 1. This trend is consistent across all datasets and missingness types, and for numerical columns as well (see Figure 6(b)). We observe similar trends for KL divergence (see Figure 6(c)), indicating that neither discrepancy nor distributional imputation quality metrics are sufficiently indicative of downstream model performance, refuting Shadbahr et al. (2023)’s claim.
@@
Imputation Fairness.
In Figure 7, we report the imputation fairness of different MVI techniques, according to F1 score difference (for categorical columns), RMSE difference (for numerical columns), and KL divergence difference (for both numerical and categorical columns, computed over the columns with nulls only) and compare it with the fairness of the downstream model, according to TPRD. We find that, while model fairness is generally agnostic to missingness type (as discussed in Section 4.2), imputation fairness is highly sensitive to missingness mechanism. For example, in Figure 7(c), miss-forest has good imputation fairness on german under MAR (KL difference of -0.4) but significant imputation unfairness under MCAR (KL difference of 2.25), MNAR (KL difference of 1.1) and mixed missingness (KL difference of 1.4).
@@ 5.1. Correctness of the Predictive Model
Test set missingness.
Figure 49 in Appendix D.2 shows the F1 of the Random Forest model on diabetes as a function of test missingness rate. We find that F1 generally decreases with an increase in test missingness, corroborating the findings of  Shadbahr et al. (2023) and Miao et al. (2022). A notable exception is miss-forest, which is robust to both forms of missingness shift such as changing missingness rates (as observed in (Miao et al., 2022)) and missingness mechanisms. As for training set missingness, F1 decreases with an increase in test missingness most steeply under MNAR (both during training and test), further supporting the findings of Miao et al. (2022). And, once again, clustering is an exception to this trend, instead showing invariance to test missingness rates under MNAR train (irrespective of test missingness) but not under MCAR and MAR train.
@@ 5.2. Fairness of the Predictive Model
Test set missingness.
We measured the impact of test missingness rate on fairness and found that fairness is highly sensitive to shifts in test missingness. Figure 50 in Appendix 49 shows the TPRD of the Random Forest model on diabetes as a function of test missingness rate. As before, fairness is highly sensitive to shifts in test missingness. For boostclean, datawig, and mnar-pvae, TPRD generally increases (fairness worsens) with an increase in test error rate, although this increase is not monotonic. miss-forest, auto-ml, and nomi are particularly robust to increases in test missingness rate under all missingness scenarios. For simpler methods such as deletion, clustering, median-mode and median-dummy, there is no consistent trend, even when missingness types remain the same and only missingness rates change between train and test. For example with deletion and clustering, TPRD increases with increase in test missingness rate in scenario S1 (MCAR train, MCAR test) and S3 (MNAR train, MNAR test), but monotonically decreases with test missingness rate in scenario S2 (MAR train, MAR test).
In summary,
our results corroborate the findings of Guha et al. (2024), and are a cause for concern as they indicate that the MVI techniques that perform best during development may not preserve fairness post-deployment, where shifts are likely to occur.
@@ 5.3. Stability of the Predictive Model
Test set missingness.
Figure 51 in Appendix 50 shows the Label Stability of the Random Forest model on diabetes as a function of test missingness rate. We find that test missingness rate has little effect on model stability under all missingness settings, with the exception of clustering, which shows lower label stability at higher missingness rates, most pronounced in scenario S7 (MAR train, MNAR test).
@@ 7. Summary of Experimental Findings
Interest in deep learning for tabular data has surged (Hwang and Song, 2023; Gorishniy et al., 2021; Borisov et al., 2022; Ye et al., 2024; Fayaz et al., 2022). Recall that Figures 3, 4, and 5 show only the best-performing models (by F1) for clarity. Tree-based methods like random forests and gradient-boosted trees outperform deep table-learning models like GANDALF (Joseph and Raj, 2022) on small datasets (diabetes, german), aligning with prior findings (Grinsztajn et al., 2022; Shwartz-Ziv and Armon, 2022). However, GANDALF excels on larger datasets (heart, folk-employment), highlighting the increasing effectiveness of deep learning across diverse data modalities.
@@ 8. Conclusions, Limitations and Future Work
Limitations.
We started with clean datasets, carefully designed meaningful missingness scenarios, and simulated them using error injectors. A limitation of our work is that we evaluate performance on synthetically generated — rather than on naturally occurring — missingness. This is a common limitation of work in this area: we often do not have a clean ground truth to compare against, and so we make the assumption that injecting synthetic errors will only exacerbate the effect of existing (unknown) errors. With this assumption, any degradation in performance we see is an underestimate of the possible degradation we can expect by comparing with an ideal clean ground truth.
Future work.
Creating research datasets with naturally occurring missingness shifts is a promising direction. While predictive performance, fairness, and stability are often treated as orthogonal, we uncover trade-offs driven by missingness, imputation choice, and, in some cases, model architecture. The absence of a universally best imputation method highlights the need for holistic metrics, evaluation procedures, and techniques grounded in the data lifecycle. Missing value imputation remains a critical aspect of responsible data engineering, to which our work makes a contribution.
@@
This supplemental material provides significantly more detail regarding the results presented in this paper. The organization is as follows: We first provide additional details on the implementation of the evaluation suite, which include a broad description and configuration of the MVI techniques used in our study (Appendix A.1) and an explanation of the corrections and enhancements made to existing benchmarks and MVI techniques to ensure the development of our truly fair evaluation suite (Appendix A.2). Next, we extend the description of our experimental settings to facilitate reproducibility (Appendix B). Specifically, Appendix B.1 provides detailed information about our experimental setup on the cluster, while Appendix B.2 extends our methodology for simulating missingness in other datasets, including proportions and base rates of protected groups, realistic missingness scenarios, feature correlation coefficients with the target, and feature importance for each dataset. In Appendix C, we offer additional results to those discussed in Section 4. In particular, Appendix C.1 presents the imputation quality of MVI techniques across all datasets from the perspectives of accuracy and training time, while Appendix C.2 showcases the fairness of predictive models measured by metrics beyond the TPRD mentioned in Section 4 and includes scatter plots for all imputers that extend scatter plots from Section 4.4. Then, Appendix D presents supplementary experimental findings on the missingness shift discussed in Section 5. Specifically, Appendix D.1 examines the accuracy, stability, and fairness dimensions of model performance under missingness shift with fixed train and test error rates, highlighting its effects on various model types and datasets. Furthermore, Appendix D.2 explores further experimental results on model performance under missingness shift with variable train and test error rates. Finally, Appendix E presents correlation plots for other missingness settings than in Section 7.
@@ C.1. Accuracy and Efficiency of MVI Techniques
C.2. Additional Accuracy and Fairness Metrics of Predictive Model
In this section, we extend the results discussed in Section 4 by presenting plots for additional popular accuracy and fairness metrics of downstream model performance. Figures 20, 21, 22, and 23 illustrate Accuracy, Disparate Impact, Selection Rate Difference, and True Negative Rate Difference of best performing models, respectively, for different imputation techniques, datasets, and missingness mechanisms. Figures 24 and 25 present scatter plots for all imputers that extend scatter plots from Section 4.4. Fairness metrics are calculated for an intersectional group when multiple sensitive attributes are present in the dataset, and for a single group when there is only one sensitive attribute (see Table 2 for reference).
@@ C.2. Additional Accuracy and Fairness Metrics of Predictive Model
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 20. Accuracy of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Datasets are ordered in increasing order by size. The blue line shows performance of a model trained on clean data.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 21. Disparate Impact (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 1 are ideal/fair. Datasets are ordered in increasing order by size. The blue line indicates the performance of a model trained on clean data.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 22. Selection Rate Difference (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 0 are ideal/fair. Datasets are ordered in increasing order by size. The blue line indicates the performance of a model trained on clean data.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
(d) Mixed missingness (MCAR & MAR & MNAR)

### v5

@@ 4. Single and Multi-Mechanism Missingness
Different models are the best-performing on different datasets. In Figures 2223 and 24, we report on the best-performing models (according to F1) for five most representative datasets per experiment, and compare performance against a model trained on clean data. Complete results are available in Appendix C.2.
@@
Figure 3. F1 of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Datasets are in increasing order by size. Blue line shows median performance of the model trained on clean data.
@@ 4.1. Correctness of the Predictive Model
Note that the F1 score on the bank dataset is extremely low (0.32). This is due to severe class imbalance in the dataset (the base rate is 0.117, as reported in Table 9 in Appendix B). Further, from Figure 25 in Appendix C.2, we can see that the accuracy of the lgbm__\__clf is 0.89, which is marginally better than the accuracy reported using Logistic Regression (0.885) in (Le Quy et al., 2022) in Table 15. Also note that a model trained on imputed data can sometimes outperform one trained on clean data, as for german and heart on MCAR. We hypothesize that this happens when the model trained on cleaned data learns spurious correlations (e.g., due to noisy or erroneous values, especially in continuous features). However, when we simulate nulls, MVI techniques that leverage information from other rows could impute more accurate values, breaking some of these spurious correlations and resulting in slightly better performance.
@@
Figure 4. True Positive Rate Difference (unfairness) of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets ($x$-axis), and missingness mechanisms (subplots). Values close to 0 are desirable. Datasets are in increasing order by size. Blue line shows median TPRD of the model trained on clean data.
@@ 4.2. Fairness of the Predictive Model
Figure 5. Label Stability of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 1 are desirable. Blue line shows median performance of the model trained on clean data.
@@ 4.3. Stability of the Predictive Model
Figure 24 shows label stability of models trained using different MVI techniques. In line with conventional wisdom (Domingos, 2000), stability depends most strongly on dataset characteristics (mainly size) and model type. Impute-then-classify approaches like miss-forest, auto-ml and datawig, which were the best-performing according to F1, also match the stability of the model trained on clean data under all missingness types and dataset sizes. nomi, which performs best on accuracy and training time, demonstrates comparable stability to these top techniques on all datasets. clustering, which as was poor-performing according to F1 (correctness) is also poor-performing according to stability on small datasets (diabetes and german, with 905 and 1k samples, respectively). A surprising exception is under mixed missingness on german, where clustering is competitive with other impute-then-classify approaches, despite not being competitive on MCAR, MAR or MNAR individually. The reason is that imputation accuracy of MVI techniques impacts data uncertainty, which then directly impacts overall model uncertainty (Gal et al., 2016).
Deletion worsens stability compared to the model trained on clean data for all missingness types on diabetes, but, notably, only under MNAR on german. boostclean, which constitutes its own model type, shows a deviation from the stability of the model trained on clean data on all datasets except bank: worsening stability compared to the clean model on folk-income, law-school and folk-employment (except under MCAR), but, surprisingly, improving it on heart, even under mixed missingness.
@@ 4.4. Imputation Quality and Fairness
Imputation Quality. We find that MVI techniques with vastly different imputation quality can give rise to models with comparable F1 scores, indicating that imputation correctness is not sufficiently predictive of model correctness (Zhang and Long, 2021a). For example, in Figure 6(a) on the diabetes dataset, median-dummy has an F1 score (imputation) close to 0; auto-ml, miss-forest and nomi have an F1 score (imputation) close to 1; and all other imputers have an F1 score (imputation) between 0.5-0.6, but the models trained with all of these techniques have an F1 close to 1. This trend is consistent across all datasets and missingness types, and for numerical columns as well (see Figure 6(b)). We observe similar trends for KL divergence (see Figure 6(c)), indicating that neither discrepancy nor distributional imputation quality metrics are sufficiently indicative of downstream model performance, refuting Shadbahr et al. (2023)’s claim.
@@
Imputation Fairness. In Figure 7, we report the imputation fairness of different MVI techniques, according to F1 score difference (for categorical columns), RMSE difference (for numerical columns), and KL divergence difference (for both numerical and categorical columns, computed over the columns with nulls only) and compare it with the fairness of the downstream model, according to TPRD. We find that, while model fairness is generally agnostic to missingness type (as discussed in Section 4.2), imputation fairness is highly sensitive to missingness mechanism. For example, in Figure 7(c), miss-forest has good imputation fairness on german under MAR (KL difference of -0.4) but significant imputation unfairness under MCAR (KL difference of 2.25), MNAR (KL difference of 1.1) and mixed missingness (KL difference of 1.4).
@@ 5.1. Correctness of the Predictive Model
Test set missingness. Figure 52 in Appendix D.2 shows the F1 of the Random Forest model on diabetes as a function of test missingness rate. We find that F1 generally decreases with an increase in test missingness, corroborating the findings of  Shadbahr et al. (2023) and Miao et al. (2022). A notable exception is miss-forest, which is robust to both forms of missingness shift such as changing missingness rates (as observed in (Miao et al., 2022)) and missingness mechanisms. As for training set missingness, F1 decreases with an increase in test missingness most steeply under MNAR (both during training and test), further supporting the findings of Miao et al. (2022). And, once again, clustering is an exception to this trend, instead showing invariance to test missingness rates under MNAR train (irrespective of test missingness) but not under MCAR and MAR train.
@@ 5.2. Fairness of the Predictive Model
Test set missingness. We measured the impact of test missingness rate on fairness and found that fairness is highly sensitive to shifts in test missingness. Figure 53 in Appendix 52 shows the TPRD of the Random Forest model on diabetes as a function of test missingness rate. As before, fairness is highly sensitive to shifts in test missingness. For boostclean, datawig, and mnar-pvae, TPRD generally increases (fairness worsens) with an increase in test error rate, although this increase is not monotonic. miss-forest, auto-ml, and nomi are particularly robust to increases in test missingness rate under all missingness scenarios. For simpler methods such as deletion, clustering, median-mode and median-dummy, there is no consistent trend, even when missingness types remain the same and only missingness rates change between train and test. For example with deletion and clustering, TPRD increases with increase in test missingness rate in scenario S1 (MCAR train, MCAR test) and S3 (MNAR train, MNAR test), but monotonically decreases with test missingness rate in scenario S2 (MAR train, MAR test).
In summary, our results corroborate the findings of Guha et al. (2024), and are a cause for concern as they indicate that the MVI techniques that perform best during development may not preserve fairness post-deployment, where shifts are likely to occur.
@@ 5.3. Stability of the Predictive Model
Test set missingness. Figure 54 in Appendix 53 shows the Label Stability of the Random Forest model on diabetes as a function of test missingness rate. We find that test missingness rate has little effect on model stability under all settings except clustering, which shows lower label stability at higher missingness rates, most pronounced in scenario S7 (MAR train, MNAR test).
@@ 7. Summary of Experimental Findings
Interest in deep learning for tabular data has surged (Hwang and Song, 2023; Gorishniy et al., 2021; Borisov et al., 2022; Ye et al., 2024; Fayaz et al., 2022). Recall that Figures 22, 23, and 24 show only the best-performing models (by F1) for clarity. Tree-based methods like random forests and gradient-boosted trees outperform deep table-learning models like GANDALF (Joseph and Raj, 2022) on small datasets (diabetes, german), aligning with prior findings (Grinsztajn et al., 2022; Shwartz-Ziv and Armon, 2022). However, GANDALF excels on larger datasets (heart, folk-employment), highlighting the increasing effectiveness of deep learning across diverse data modalities.
@@ 8. Conclusions, Limitations and Future Work
Limitations. We started with clean datasets, carefully designed meaningful missingness scenarios, and simulated them using error injectors. A limitation of our work is that we evaluate performance on synthetically generated — rather than on naturally occurring — missingness. This is a common limitation of work in this area: we often do not have a clean ground truth to compare against, and so we make the assumption that injecting synthetic errors will only exacerbate the effect of existing (unknown) errors. With this assumption, any degradation in performance we see is an underestimate of the possible degradation we can expect by comparing with an ideal clean ground truth.
Future work. Creating research datasets with naturally occurring missingness shifts is a promising direction. While predictive performance, fairness, and stability are often treated as orthogonal, we uncover trade-offs driven by missingness, imputation choice, and, in some cases, model architecture. The absence of a universally best imputation method highlights the need for holistic metrics, evaluation procedures, and techniques grounded in the data lifecycle. Missing value imputation remains a critical aspect of responsible data engineering, to which our work makes a contribution.
@@
This supplemental material provides significantly more detail regarding the results presented in this paper. The organization is as follows: We first provide additional details on the implementation of the evaluation suite, which include a broad description and configuration of the MVI techniques used in our study (Appendix A.1) and an explanation of the corrections and enhancements made to existing benchmarks and MVI techniques to ensure the development of our truly fair evaluation suite (Appendix A.2). Next, we extend the description of our experimental settings to facilitate reproducibility (Appendix B). Specifically, Appendix B.1 provides detailed information about our experimental setup on the cluster, while Appendix B.2 extends our methodology for simulating missingness in other datasets, including proportions and base rates of protected groups, realistic missingness scenarios, feature correlation coefficients with the target, and feature importance for each dataset. In Appendix C, we offer additional results to those discussed in Section 4. In particular, Appendix C.1 presents the imputation quality of MVI techniques across all datasets from the perspectives of accuracy and training time, while Appendix C.2 showcases additional metrics for predictive model performance beyond those mentioned in Section 4 and includes scatter plots for all imputers that extend scatter plots from Section 4.4. Then, Appendix D presents supplementary experimental findings on the missingness shift discussed in Section 5. Specifically, Appendix D.1 examines the accuracy, stability, and fairness dimensions of model performance under missingness shift with fixed train and test error rates, highlighting its effects on various model types and datasets. Furthermore, Appendix D.2 explores further experimental results on model performance under missingness shift with variable train and test error rates. Finally, Appendix E presents correlation plots for other missingness settings than in Section 7.
@@ C.1. Accuracy and Efficiency of MVI Techniques
C.2. Additional Experimental Results for Predictive Model Performance
In this section, we extend the results discussed in Section 4 by presenting plots for additional metrics of downstream model performance. Figures 22, 23, 24, 25, 26, 27, and 28 illustrate F1, True Positive Rate Difference, Label Stability, Accuracy, Disparate Impact, Selection Rate Difference, and True Negative Rate Difference of best performing models, respectively, for different imputation techniques, datasets, and missingness mechanisms. Figures 20 and 21 present scatter plots for all imputers that extend scatter plots from Section 4.4. Fairness metrics are calculated for an intersectional group when multiple sensitive attributes are present in the dataset, and for a single group when there is only one sensitive attribute (see Table 2 for reference).
@@ C.2. Additional Experimental Results for Predictive Model Performance
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 22. F1 of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Datasets are in increasing order by size. Blue line shows median performance of the model trained on clean data.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 23. True Positive Rate Difference (unfairness) of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets ($x$-axis), and missingness mechanisms (subplots). Values close to 0 are desirable. Datasets are in increasing order by size. Blue line shows median TPRD of the model trained on clean data.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 24. Label Stability of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 1 are desirable. Datasets are ordered in increasing order by size. The blue line indicates median performance of the model trained on clean data.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 25. Accuracy of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Datasets are ordered in increasing order by size. The blue line shows median performance of a model trained on clean data.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 26. Disparate Impact (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 1 are ideal/fair. Datasets are ordered in increasing order by size. The blue line indicates the median performance of a model trained on clean data.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 27. Selection Rate Difference (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 0 are ideal/fair. Datasets are ordered in increasing order by size. The blue line indicates the median performance of a model trained on clean data.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
(d) Mixed missingness (MCAR & MAR & MNAR)